OpenStudy (anonymous):

find the end behavior f(x)=(x-3)(1-x)(3x-5) Lim f(x) lim f(x) x--> infinity x--> negative infinity

OpenStudy (anonymous):

After expanding \[ f(x)=(1-x) (x-3) (3 x-5)\\ f(x)=-3 x^3+17 x^2-29 x+15\\ \]

OpenStudy (anonymous):

A polynomial near Infinity behaves like its term of highest degree, meaning like \[- 3 x^3 \] Hence it approaches -Infinity when x approaches Infinity and it approaches xxx when x approaches -Infinity. Can you find what xxx is?

OpenStudy (anonymous):

xxx?

OpenStudy (anonymous):

It means find the limit of f(x) when x approaches - Infinity

OpenStudy (anonymous):

ok so the answer would be infinity?

OpenStudy (anonymous):

Yes when x approaches - Infinty

OpenStudy (anonymous):

ok how did you find that? im really confused

OpenStudy (anonymous):

How much is \[ - 4 x^3 \] When x =-10, x=-100, x = - 1000?

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Did you understand what is going on?